# Horizon Brightened Acceleration Radiation from Massive Vector Fields

### v1

@@ II.3 Frequency labels and asymptotic kinematics
$k_{\infty}(\nu)=\sqrt{\nu^{2}-m_{V}^{2}},\qquad v_{g}(\nu)=\frac{k_{\infty}(\nu)}{\nu},$ (12)
@@
$\nu_{\text{stat}}(r)=\frac{\nu}{\sqrt{f(r)}},\qquad f(r)=1-\frac{1}{r},$ (13)
@@
By contrast, the infalling atom samples the comoving frequency given by the Doppler-gravitational invariant Carroll_2019
$\Omega(\tau,\nu)=-\,k_{\mu}u^{\mu}=\nu\!\left(\frac{dt}{d\tau}-\frac{dr_{*}}{d\tau}\right),$ (14)
where $u^{\mu}$ is the atom’s four-velocity and we have used the outgoing WKB phase $\sim e^{-i\nu(t-r_{*})}$ near the horizon. We choose the associated covector $k_{\mu}=-\partial_{\mu}(t-r_{*})=(-1,\,+\partial_{r}r_{*},\,0,\,0)=\left(-1,\,(1-1/r)^{-1},\,0,\,0\right)$, so that $k_{t}=-\nu$, $k_{r}=\nu\,dr_{*}/dr$. Using the geodesic components $u^{t}=r/(r-1)$ and $u^{r}=-1/\sqrt{r}$ from (5), the invariant contraction (14) yields
$\Omega(\tau,\nu)\equiv-\,k_{\mu}u^{\mu}=\nu\!\left(\frac{dt}{d\tau}-\frac{dr_{*}}{d\tau}\right)=\nu\!\left(\frac{r}{r-1}+\frac{\sqrt{r}}{r-1}\right)=\nu\,\frac{r+\sqrt{r}}{r-1}\,$, (15)
@@
The kinematic origin of the excitation is visualized in Fig. 1, which plots the comoving frequency ratio $\Omega(\tau,\nu)/\nu$ as a function of the radial distance. As the atom approaches the Schwarzschild radius ($r\to 1^{+}$), the redshift-Doppler factor grows without bound. This divergence or horizon brightening implies that the infalling atom perceives the finite-frequency vacuum modes of the laboratory frame as having arbitrarily high energies. Consequently, the atom can undergo transitions mediated by the counter-rotating terms in the interaction Hamiltonian, which are energetically forbidden in inertial frames but become dominant near the horizon.
@@
$\Pi_{\mu\nu}(\nu,m_{V})\equiv\sum_{\lambda}\varepsilon^{(\lambda)}_{\mu}(\nu)\,\varepsilon^{(\lambda)}_{\nu}(\nu),$ (16)
so that matrix elements for the charged-monopole and dipole detector models reduce to scalar contractions with $u^{\mu}$ or the atom’s instantaneous spatial direction, respectively. The kinematic content of this subsection is thus: (i) a hard threshold at $\nu=m_{V}$, (ii) an asymptotic phase space governed by $k_{\infty}(\nu)$, and (iii) a large, universal blue-shift of the comoving frequency $\Omega(\tau,\nu)$ dictated by (15). These ingredients will control the prefactors of the excitation probabilities, while the Planckian detailed-balance factor will originate from the phase structure implicit in (14)-(15).
@@ III Proca field on Schwarzschild
We formulate a massive spin-1 field on the Schwarzschild exterior as a Proca vector $A_{\mu}$ with field strength $F_{\mu\nu}$. The key structural features are: (i) a mass term that removes gauge redundancy and enforces the covariant Lorenz constraint $\nabla_{\mu}A^{\mu}=0$; (ii) equations of motion that reduce, in Ricci-flat backgrounds, to a vector wave equation with a mass term; and (iii) three physical polarizations. These ingredients will permit a clean separation into axial/polar sectors and a near-horizon WKB analysis in later subsections, where the outgoing mode phase aligns with the convention fixed in Section II, $e^{-i\nu(t-r_{*})}$.
@@ III.1 Equations and constraints
We work on the exterior Schwarzschild geometry already defined in Section II. The Proca action is Rosa:2011my
$S[A]=\int\!d^{4}x\,\sqrt{-g}\left(-\frac{1}{4}\,F_{\mu\nu}F^{\mu\nu}+\frac{m_{V}^{2}}{2}\,A_{\mu}A^{\mu}\right),\qquad F_{\mu\nu}\equiv\nabla_{\mu}A_{\nu}-\nabla_{\nu}A_{\mu}.$ (17)
Variation with respect to $A_{\mu}$ yields the Proca equations of motion Rosa:2011my
$\nabla_{\nu}F^{\mu\nu}+m_{V}^{2}A^{\mu}=0.$ (18)
Taking the covariant divergence of (18) gives the Proca constraint
$m_{V}^{2}\,\nabla_{\mu}A^{\mu}=0\quad\Longrightarrow\quad\;\nabla_{\mu}A^{\mu}=0\;,$ (19)
@@
It is often convenient to eliminate $F_{\mu\nu}$ using (19). Acting with $\nabla_{\nu}$ on the definition of $F_{\mu\nu}$ and using the commutator of covariant derivatives, one obtains the equivalent second-order vector equation Birrell:1982ix; Carroll_2019
$\left(\Box\,\delta^{\mu}_{\ \nu}-R^{\mu}_{\ \nu}+m_{V}^{2}\,\delta^{\mu}_{\ \nu}\right)A^{\nu}=0,\qquad\Box\equiv\nabla_{\alpha}\nabla^{\alpha}.$ (20)
In the Schwarzschild exterior $R_{\mu\nu}=0$, so (20) simplifies to
$\left(\Box+m_{V}^{2}\right)A^{\mu}=0\quad\text{with}\quad\nabla_{\mu}A^{\mu}=0.$ (21)
Equations (18)-(21) show that there are three propagating degrees of freedom, corresponding to the massive vector’s physical polarizations. For later use, we also record the symmetric stress-energy tensor obtained from (17),
$T_{\mu\nu}=F_{\mu\alpha}F_{\nu}^{\ \alpha}-\frac{1}{4}\,g_{\mu\nu}F_{\alpha\beta}F^{\alpha\beta}+m_{V}^{2}\!\left(A_{\mu}A_{\nu}-\frac{1}{2}g_{\mu\nu}A_{\alpha}A^{\alpha}\right),$ (22)
which is conserved, $\nabla^{\mu}T_{\mu\nu}=0$, when (18) holds Birrell:1982ix.
@@ III.2 Master radial equations and effective potentials
We separate the Proca field $A_{\mu}$ on the Schwarzschild exterior (with f(r)=11/rf(r)=1-1/r and tortoise $dr_{*}/dr=f^{-1}$) into vector spherical harmonics. The decomposition splits into an axial (odd-parity) sector with one master variable and a polar (even-parity) sector carrying two coupled master variables. Throughout we use the positive-frequency convention for outgoing modes fixed in Section II, namely phases $e^{-i\nu(t-r_{*})}$.
@@
$\frac{d^{2}\Psi_{\ell}^{\text{(ax)}}}{dr_{*}^{2}}+\left[\nu^{2}-V_{\ell}^{\text{(ax)}}(r)\right]\Psi_{\ell}^{\text{(ax)}}=0,$ (23)
@@
$V_{\ell}^{\text{(ax)}}(r)=f(r)\!\left[\frac{\ell(\ell+1)}{r^{2}}+m_{V}^{2}\right].\;$ (24)
@@
$\frac{d^{2}}{dr_{*}^{2}}\begin{pmatrix}\Phi_{\ell}^{(1)}\\ \Phi_{\ell}^{(2)}\end{pmatrix}+\left[\nu^{2}\mathbf{I}-\mathbf{U}_{\ell}(r)\right]\begin{pmatrix}\Phi_{\ell}^{(1)}\\ \Phi_{\ell}^{(2)}\end{pmatrix}=0,$ (25)
@@
$\mathbf{U}_{\ell}(r)=f(r)\!\begin{pmatrix}\displaystyle m_{V}^{2}+\frac{\ell(\ell+1)}{r^{2}}+\frac{\alpha_{\ell}(r)}{r^{3}}&\displaystyle\frac{\beta_{\ell}(r)}{r^{2}}\\ \displaystyle\frac{\gamma_{\ell}(r)}{r^{2}}&\displaystyle m_{V}^{2}+\frac{\ell(\ell+1)}{r^{2}}+\frac{\delta_{\ell}(r)}{r^{3}}\end{pmatrix}$. (26)
The functions $\alpha_{\ell},\beta_{\ell},\gamma_{\ell},\delta_{\ell}$ are rational in $r$ (dimensionless in our units) and arise from curvature couplings and the Proca constraint; they fall at least as $r^{-1}$. Two facts suffice for our purposes: (1) Since $\mathbf{U}_{\ell}\propto f(r)$, each entry vanishes linearly as $r\to 1^{+}$. Thus, both polar master fields share the same plane-wave behavior as the axial field:
$\Phi_{\ell}^{(i)}\sim e^{-i\nu(t-r_{*})},\qquad r\to 1^{+},\ i=1,2.$ (27)
@@
$\frac{d^{2}\Psi_{\ell}^{\text{(pol)}\pm}}{dr_{*}^{2}}+\left[\nu^{2}-V_{\ell}^{\text{(pol)}\pm}(r)\right]\Psi_{\ell}^{\text{(pol)}\pm}=0,$ (28)
with effective potentials $V_{\ell}^{\text{(pol)}\pm}(r)$ that share the same limiting behaviors as in (24), viz.
$V_{\ell}^{\text{(pol)}\pm}(r)=f(r)\!\left[m_{V}^{2}+\frac{\ell(\ell+1)}{r^{2}}\right]+\mathcal{O}\!\left(\frac{f(r)}{r^{3}}\right).$ (29)
@@
$\mathcal{W}=\Psi^{\ast}\frac{d\Psi}{dr_{*}}-\frac{d\Psi^{\ast}}{dr_{*}}\Psi$ (30)
is constant in $r_{*}$. We normalize solutions so that a unit outgoing flux at infinity corresponds to $\Psi\sim e^{-ik_{\infty}r_{*}}$ with $k_{\infty}=\sqrt{\nu^{2}-m_{V}^{2}}$. Near the horizon the same mode behaves as $e^{-i\nu(t-r_{*})}$, consistent with Section II. These conventions set the stage for the stationary-phase worldline integrals and for introducing polarization sums and greybody factors in later sections.
@@ III.3 Mode normalization and density of states
$A_{\mu}(x)=\sum_{\ell m}\sum_{\lambda=1}^{3}\int_{m_{V}}^{\infty}\!d\nu\;\left[a_{\nu\ell m\lambda}\,\mathcal{A}^{(\lambda)}_{\mu;\nu\ell m}(x)+a_{\nu\ell m\lambda}^{\dagger}\,\mathcal{A}^{(\lambda)\,*}_{\mu;\nu\ell m}(x)\right],$ (31)
@@
$\mathcal{A}^{(\lambda)}_{\mu;\nu\ell m}(x)\sim\frac{1}{r}\frac{\varepsilon^{(\lambda)}_{\mu}(\nu,\hat{\mathbf{k}})}{\sqrt{k_{\infty}(\nu)}}Y_{\ell m}(\theta,\phi)\,e^{-i\nu t}e^{-ik_{\infty}(\nu)r_{*}},\qquad k_{\infty}(\nu)=\sqrt{\nu^{2}-m_{V}^{2}}.$ (32)
@@
$\sum_{\lambda=1}^{3}\varepsilon^{(\lambda)}_{\mu}\,\varepsilon^{(\lambda)}_{\nu}=-\eta_{\mu\nu}+\frac{k_{\mu}k_{\nu}}{m_{V}^{2}},$ (33)
@@
To connect with the cavity description in Section II, we place reflecting walls at tortoise locations $r_{*}^{(b)}<r_{*}<r_{*}^{(t)}$ of length $L_{*}\equiv r_{*}^{(t)}-r_{*}^{(b)}$. Standing-wave conditions quantize the asymptotic wavenumber as $k_{\infty}L_{*}=\pi n$ (integer $n$). In the continuum limit $L_{*}\to\infty$ with a narrow frequency window, the number of modes in $\nu,\nu+d\nu]$ for a fixed $(\ell,m,\lambda)$ is
$\rho(\nu)\,d\nu=\frac{L_{*}}{\pi}\,\frac{dk_{\infty}}{d\nu}\,d\nu=\frac{L_{*}}{\pi}\,\frac{\nu}{\sqrt{\nu^{2}-m_{V}^{2}}}\,\Theta(\nu-m_{V})\,d\nu\;,$ (34)
@@
Near the horizon, all master variables in Section III.2 reduce to plane waves with phase $e^{-i\nu(t-r_{*})}$; the normalization at infinity then propagates inward by constancy of the Wronskian. These conventions ensure that transition amplitudes derived along the infalling worldline carry the correct polarization sum (33) and the correct kinematic weight through $k_{\infty}(\nu)$ and $\rho(\nu)$ in (34).

### v2

@@ II.3 Frequency labels and asymptotic kinematics
$k_{\infty}(\nu)=\sqrt{\nu^{2}-m_{V}^{2}},\qquad v_{g}(\nu)=\frac{k_{\infty}(\nu)}{\nu},$ (II.12)
@@
$\nu_{\text{stat}}(r)=\frac{\nu}{\sqrt{f(r)}},\qquad f(r)=1-\frac{1}{r},$ (II.13)
@@
By contrast, the infalling atom samples the comoving frequency given by the Doppler-gravitational invariant [17]
$\Omega(\tau,\nu)=-\,k_{\mu}u^{\mu}=\nu\!\left(\frac{dt}{d\tau}-\frac{dr_{*}}{d\tau}\right),$ (II.14)
where $u^{\mu}$ is the atom’s four-velocity and we have used the outgoing WKB phase $\sim e^{-i\nu(t-r_{*})}$ near the horizon. We choose the associated covector $k_{\mu}=-\partial_{\mu}(t-r_{*})=(-1,\,+\partial_{r}r_{*},\,0,\,0)=\left(-1,\,(1-1/r)^{-1},\,0,\,0\right)$, so that $k_{t}=-\nu$, $k_{r}=\nu\,dr_{*}/dr$. Using the geodesic components $u^{t}=r/(r-1)$ and $u^{r}=-1/\sqrt{r}$ from (II.5), the invariant contraction (II.14) yields
$\Omega(\tau,\nu)$ $\equiv-\,k_{\mu}u^{\mu}=\nu\!\left(\frac{dt}{d\tau}-\frac{dr_{*}}{d\tau}\right)$
$=\nu\!\left(\frac{r}{r-1}+\frac{\sqrt{r}}{r-1}\right)=\nu\,\frac{r+\sqrt{r}}{r-1}\,$, (II.15)
@@
The kinematic origin of the excitation is visualized in Fig. 1, which plots the comoving frequency ratio $\Omega(\tau,\nu)/\nu$ as a function of the radial distance. As the atom approaches the Schwarzschild radius ($r\to 1^{+}$), the redshift-Doppler factor grows without bound. This divergence or horizon brightening implies that the infalling atom perceives the finite-frequency vacuum modes of the laboratory frame as having arbitrarily high energies. Consequently, the atom can undergo transitions mediated by the counterrotating terms in the interaction Hamiltonian, which are energetically forbidden in inertial frames but become dominant near the horizon.
@@
$\Pi_{\mu\nu}(\nu,m_{V})\equiv\sum_{\lambda}\varepsilon^{(\lambda)}_{\mu}(\nu)\,\varepsilon^{(\lambda)}_{\nu}(\nu),$ (II.16)
so that matrix elements for the charged-monopole and dipole detector models reduce to scalar contractions with $u^{\mu}$ or the atom’s instantaneous spatial direction, respectively. The kinematic content of this subsection is thus: (i) a hard threshold at $\nu=m_{V}$, (ii) an asymptotic phase space governed by $k_{\infty}(\nu)$, and (iii) a large, universal blueshift of the comoving frequency $\Omega(\tau,\nu)$ dictated by (II.15). These ingredients will control the prefactors of the excitation probabilities, while the Planckian detailed-balance factor will originate from the phase structure implicit in (II.14)–(II.15).
@@ III Proca field on Schwarzschild
We formulate a massive spin-1 field on the Schwarzschild exterior as a Proca vector $A_{\mu}$ with field strength $F_{\mu\nu}$. The key structural features are: (i) a mass term that removes gauge redundancy and enforces the covariant Lorenz constraint $\nabla_{\mu}A^{\mu}=0$; (ii) equations of motion that reduce, in Ricci-flat backgrounds, to a vector wave equation with a mass term; and (iii) three physical polarizations. These ingredients will permit a clean separation into axial/polar sectors and a near-horizon WKB analysis in later subSecs., where the outgoing mode phase aligns with the convention fixed in Sec. II, $e^{-i\nu(t-r_{*})}$.
@@ III.1 Equations and constraints
We work on the exterior Schwarzschild geometry already defined in Sec. II. The Proca action is [41]
$S[A]$ $=\int\!d^{4}x\,\sqrt{-g}\left(-\frac{1}{4}\,F_{\mu\nu}F^{\mu\nu}+\frac{m_{V}^{2}}{2}\,A_{\mu}A^{\mu}\right),$
$F_{\mu\nu}$ $\equiv\nabla_{\mu}A_{\nu}-\nabla_{\nu}A_{\mu}.$ (III.1)
Variation with respect to $A_{\mu}$ yields the Proca equations of motion [41]
$\nabla_{\nu}F^{\mu\nu}+m_{V}^{2}A^{\mu}=0.$ (III.2)
Taking the covariant divergence of (III.2) gives the Proca constraint
$m_{V}^{2}\,\nabla_{\mu}A^{\mu}=0\quad\Longrightarrow\quad\;\nabla_{\mu}A^{\mu}=0\;,$ (III.3)
@@
It is often convenient to eliminate $F_{\mu\nu}$ using (III.3). Acting with $\nabla_{\nu}$ on the definition of $F_{\mu\nu}$ and using the commutator of covariant derivatives, one obtains the equivalent second-order vector equation [9, 17]
$\left(\Box\,\delta^{\mu}_{\ \nu}-R^{\mu}_{\ \nu}+m_{V}^{2}\,\delta^{\mu}_{\ \nu}\right)A^{\nu}=0,\qquad\Box\equiv\nabla_{\alpha}\nabla^{\alpha}.$ (III.4)
In the Schwarzschild exterior $R_{\mu\nu}=0$, so (III.4) simplifies to
$\left(\Box+m_{V}^{2}\right)A^{\mu}=0\quad\text{with}\quad\nabla_{\mu}A^{\mu}=0.$ (III.5)
Equations (III.2)–(III.5) show that there are three propagating degrees of freedom, corresponding to the massive vector’s physical polarizations. For later use, we also record the symmetric stress-energy tensor obtained from (III.1),
$T_{\mu\nu}$ $=F_{\mu\alpha}F_{\nu}^{\ \alpha}-\frac{1}{4}\,g_{\mu\nu}F_{\alpha\beta}F^{\alpha\beta}+m_{V}^{2}$
$\times\!\left(A_{\mu}A_{\nu}-\frac{1}{2}g_{\mu\nu}A_{\alpha}A^{\alpha}\right),$ (III.6)
which is conserved, $\nabla^{\mu}T_{\mu\nu}=0$, when (III.2) holds [9].
@@ III.2 Master radial equations and effective potentials
We separate the Proca field $A_{\mu}$ on the Schwarzschild exterior [with f(r)=11/rf(r)=1-1/r and tortoise $dr_{*}/dr=f^{-1}$] into vector spherical harmonics. The decomposition splits into an axial (odd-parity) sector with one master variable and a polar (even-parity) sector carrying two coupled master variables. Throughout we use the positive-frequency convention for outgoing modes fixed in Sec. II, namely phases $e^{-i\nu(t-r_{*})}$.
@@
$\frac{d^{2}\Psi_{\ell}^{\text{(ax)}}}{dr_{*}^{2}}+\left[\nu^{2}-V_{\ell}^{\text{(ax)}}(r)\right]\Psi_{\ell}^{\text{(ax)}}=0,$ (III.7)
@@
$V_{\ell}^{\text{(ax)}}(r)=f(r)\!\left[\frac{\ell(\ell+1)}{r^{2}}+m_{V}^{2}\right].\;$ (III.8)
@@
$\frac{d^{2}}{dr_{*}^{2}}\begin{pmatrix}\Phi_{\ell}^{(1)}\\ \Phi_{\ell}^{(2)}\end{pmatrix}+\left[\nu^{2}\mathbf{I}-\mathbf{U}_{\ell}(r)\right]\begin{pmatrix}\Phi_{\ell}^{(1)}\\ \Phi_{\ell}^{(2)}\end{pmatrix}=0,$ (III.9)
@@
$\mathbf{U}_{\ell}(r)=f(r)\!\begin{pmatrix}\displaystyle m_{V}^{2}+\frac{\ell(\ell+1)}{r^{2}}+\frac{\alpha_{\ell}(r)}{r^{3}}&\displaystyle\frac{\beta_{\ell}(r)}{r^{2}}\\ \displaystyle\frac{\gamma_{\ell}(r)}{r^{2}}&\displaystyle m_{V}^{2}+\frac{\ell(\ell+1)}{r^{2}}+\frac{\delta_{\ell}(r)}{r^{3}}\end{pmatrix}$. (III.10)
The functions $\alpha_{\ell},\beta_{\ell},\gamma_{\ell},\delta_{\ell}$ are rational in $r$ (dimensionless in our units) and arise from curvature couplings and the Proca constraint; they fall at least as $r^{-1}$. Two facts suffice for our purposes: (1) Since $\mathbf{U}_{\ell}\propto f(r)$, each entry vanishes linearly as $r\to 1^{+}$. Thus, both polar master fields share the same plane wave behavior as the axial field:
$\Phi_{\ell}^{(i)}\sim e^{-i\nu(t-r_{*})},\qquad r\to 1^{+},\ i=1,2.$ (III.11)
@@
$\frac{d^{2}\Psi_{\ell}^{\text{(pol)}\pm}}{dr_{*}^{2}}+\left[\nu^{2}-V_{\ell}^{\text{(pol)}\pm}(r)\right]\Psi_{\ell}^{\text{(pol)}\pm}=0,$ (III.12)
with effective potentials $V_{\ell}^{\text{(pol)}\pm}(r)$ that share the same limiting behaviors as in (III.8), viz.
$V_{\ell}^{\text{(pol)}\pm}(r)=f(r)\!\left[m_{V}^{2}+\frac{\ell(\ell+1)}{r^{2}}\right]+\mathcal{O}\!\left(\frac{f(r)}{r^{3}}\right).$ (III.13)
@@
$\mathcal{W}=\Psi^{\ast}\frac{d\Psi}{dr_{*}}-\frac{d\Psi^{\ast}}{dr_{*}}\Psi$ (III.14)
is constant in $r_{*}$. We normalize solutions so that a unit outgoing flux at infinity corresponds to $\Psi\sim e^{-ik_{\infty}r_{*}}$ with $k_{\infty}=\sqrt{\nu^{2}-m_{V}^{2}}$. Near the horizon the same mode behaves as $e^{-i\nu(t-r_{*})}$, consistent with Sec. II. These conventions set the stage for the stationary-phase worldline integrals and for introducing polarization sums and greybody factors in later Secs..
@@ III.3 Mode normalization and density of states
$A_{\mu}(x)$ $=\sum_{\ell m}\sum_{\lambda=1}^{3}\int_{m_{V}}^{\infty}\!d\nu\;\Big[a_{\nu\ell m\lambda}\,\mathcal{A}^{(\lambda)}_{\mu;\nu\ell m}(x)$
$+a_{\nu\ell m\lambda}^{\dagger}\,\mathcal{A}^{(\lambda)\,*}_{\mu;\nu\ell m}(x)\Big],$ (III.15)
@@
$\mathcal{A}^{(\lambda)}_{\mu;\nu\ell m}(x)$ $\sim\frac{1}{r}\frac{\varepsilon^{(\lambda)}_{\mu}(\nu,\hat{\mathbf{k}})}{\sqrt{k_{\infty}(\nu)}}Y_{\ell m}(\theta,\phi)\,e^{-i\nu t}e^{-ik_{\infty}(\nu)r_{*}},$
$k_{\infty}(\nu)$ $=\sqrt{\nu^{2}-m_{V}^{2}}.$ (III.16)
@@
$\sum_{\lambda=1}^{3}\varepsilon^{(\lambda)}_{\mu}\,\varepsilon^{(\lambda)}_{\nu}=-\eta_{\mu\nu}+\frac{k_{\mu}k_{\nu}}{m_{V}^{2}},$ (III.17)
@@
To connect with the cavity description in Sec. II, we place reflecting walls at tortoise locations $r_{*}^{(b)}<r_{*}<r_{*}^{(t)}$ of length $L_{*}\equiv r_{*}^{(t)}-r_{*}^{(b)}$. Standing-wave conditions quantize the asymptotic wave number as $k_{\infty}L_{*}=\pi n$ (integer $n$). In the continuum limit $L_{*}\to\infty$ with a narrow frequency window, the number of modes in $\nu,\nu+d\nu]$ for a fixed $(\ell,m,\lambda)$ is
$\rho(\nu)\,d\nu=\frac{L_{*}}{\pi}\,\frac{dk_{\infty}}{d\nu}\,d\nu=\frac{L_{*}}{\pi}\,\frac{\nu}{\sqrt{\nu^{2}-m_{V}^{2}}}\,\Theta(\nu-m_{V})\,d\nu\;,$ (III.18)
@@
Near the horizon, all master variables in Sec. III.2 reduce to plane waves with phase $e^{-i\nu(t-r_{*})}$; the normalization at infinity then propagates inward by constancy of the Wronskian. These conventions ensure that transition amplitudes derived along the infalling worldline carry the correct polarization sum (III.17) and the correct kinematic weight through $k_{\infty}(\nu)$ and $\rho(\nu)$ in (III.18).